OpenStudy (anonymous):

A rocket is shot into the air from the ground so that its height H in feet after t seconds is given by H(t) = 200t + 10t^2 − 4t^3. Find the maximum and minimum velocity of the rocket over the interval on which it travels upward. Find the maximum and minimum velocity of the rocket over the first second of its journey.

OpenStudy (anonymous):

do you know how to use derivatives?

OpenStudy (anonymous):

are you in calculus? what subject?

OpenStudy (anonymous):

@ kimmy0394

OpenStudy (anonymous):

can you do derivative of H, d(H)dt ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok try that and lets see what you get on dH/dt

OpenStudy (anonymous):

f'(x)= 200+20t-12t^2

OpenStudy (anonymous):

f''(x)= 20-24t

OpenStudy (anonymous):

ok good, that dH/dt = velocity v for maximum H ,, equate dH/dt=0, and find t, see what you are getting

OpenStudy (anonymous):

because max H is attained when v=dH/dt=0

OpenStudy (anonymous):

find t there then sub it to H(t) to get max H

OpenStudy (anonymous):

so far you did velocity v=dH/dt=f'(x)= 200+20t-12t^2 now equate it to ZERO,and find t, that is velocity v=dH/dt=f'(x)= 200+20t-12t^2=0 or 200+20t-12t^2=0, find t?

OpenStudy (anonymous):

hint: use quadratic formula here

OpenStudy (anonymous):

\[t=x=\frac{ -b+-\sqrt{b^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

a=-12 , b=20, c=200 plug them to your quadratic formula and fint t or x

OpenStudy (anonymous):

-20+/- sqareroot (-19200) / 24

OpenStudy (anonymous):

do the + first then the -

OpenStudy (anonymous):

t=(-b+sqrt(b^2 -4ac))/2a then do the negative t=(-b-sqrt(b^2 -4ac))/2a

OpenStudy (anonymous):

watch carefullt you signs specially on negative sign

OpenStudy (anonymous):

\[t=\frac{ -20+\sqrt{29^{2}-4(-12)(200)} }{ 2(-12) }\]

OpenStudy (anonymous):

29^2? or 20^2?

OpenStudy (anonymous):

\[t=\frac{ -20+100 }{ -24 }\]

OpenStudy (anonymous):

and \[t=\frac{ -20-100 }{ -24 }\]

OpenStudy (anonymous):

so whats the t1 and t2?

OpenStudy (anonymous):

-80/24 and 120/24

OpenStudy (anonymous):

yes do it in whole number or mixed number...

OpenStudy (anonymous):

so how can you determine max and min

OpenStudy (anonymous):

ok good t1=-3.33 sec forget this negative value and t2= 5 sec ...plug this into you HJ that will be your maxH

OpenStudy (anonymous):

ok now that we know how to find the maxH from the question now, can you find the MAX V?

OpenStudy (anonymous):

you plug in 5 to derivative

OpenStudy (anonymous):

t= 5sec is to find your maxH, if you have velocity V= 200+20t-12t^2 can you try to find the maxV same as finding maxH

OpenStudy (anonymous):

hint find the derivative of V, then equate it to zero and find t then sub it to V to get the maxV or minV

OpenStudy (anonymous):

dV/dt=? then equate it to zero dV/dt=0=?

OpenStudy (anonymous):

try it first and let me see what u get

OpenStudy (anonymous):

from f'(x)= 200+20t-12t^2 =velocity V V= 200+20t-12t^2 now do dV/dt=0, fint t? to find the max and min V

OpenStudy (anonymous):

from V= 200+20t-12t^2 is dV/dt=20 - 24t ?

OpenStudy (anonymous):

if yes, find dV/dt=20 -24 t =0, find t=?

OpenStudy (anonymous):

is t=20/24 =0.833 sec ? if yes, sub it to V= 200+20t-12t^2 V= 200 +20(0.833) - 12(0.833)^2 max V=?

OpenStudy (anonymous):

you can use calculator here... lol...:D

OpenStudy (anonymous):

is max V=208.33 m/sec ?

OpenStudy (anonymous):

my pc froze hold on a sec

OpenStudy (anonymous):

contin ue solving them ..lol :D

OpenStudy (anonymous):

check that out..you can use wolfram on all of your work, for graphing etc

OpenStudy (anonymous):

so theres no min value of V here only maxV

OpenStudy (anonymous):

Find the maximum and minimum velocity of the rocket over the first second of its journey.

OpenStudy (anonymous):

dH/dt= f'(t)=V= 200+20t-12t^2 for the ist sec, t=1 sec sub this to V= 200+20t-12t^2

OpenStudy (anonymous):

is it for t=1 sec... max V= 200+20(1)-12(1)^2 = 208 ?